OpenStudy (anonymous):

Find the midpoint between the two points. (3, 0) (5, 4) a. (3, 4) b. (1, 2) c. (4, 2) d. (0, 5)

OpenStudy (anonymous):

c\

OpenStudy (dumbcow):

midpoint is really easy...just take the average of x_coordinates and avg of y_coordinates

OpenStudy (anonymous):

(3+5)/2=4 (0=4)/2=2

OpenStudy (anonymous):

0+4 *

OpenStudy (amistre64):

yep; add em and divide by 2 (3, 0) (5, 4) ----- (8, 4)/2 = (4, 2)

OpenStudy (anonymous):

the answer is C using midpoint formula (8/2;4/2) =(4;2)